# Distributional Diffusion Models with Scoring Rules

### v2

@@ F.2 Connection with DDIM
Note that in (Song et al., 2021a), we $\alpha_{t}^{2}$ is replaced with $\alpha_{t}$ and $\eta_{s,t}$ is denoted $\sigma_{t}$. Recall that in our introduction of diffusion models in Section 2.1, see (6) we have
$\mu_{s,t}(x_{0},x_{t})=\varepsilon^{2}r_{1,2}(s,t)+(1-\varepsilon^{2})r_{0,1})x_{t}$ (180)
@@ I.2 Image space experiments
For all the models, we sweep over weighting $w_{t}$ (see (8), (14)), where we either do not use any weighting or we use sigmoid weighting, see (Kingma et al., 2021), weighting with bias parameters $\{-2,-1,0,1,2\}$. When we train classical diffusion model, we always use velocity as a prediction target, i.e., the target is $X_{0}-X_{1}$. When we train distributional model, we use $X_{0}$ as prediction target. For distributional models, we additionally sweep over $\lambda\in\{0,0.1,0.5,1.0\}$, $\beta\in\{0.0001,0.001,0.01,0.1,0.5,1.0,1.5,1.99,2.0\}$. When we use kernel diffusion losses, see  Section 2.2 and Section 3, for (imq) kernel (9a), we sweep over $c\in\{0.01,0.1,0.3,0.5,1.0,1.5,2.0,10.0,50.0,100.0\}$. For for (rbf) kernel (9b) and for (exp) kernel (9c), we firstly compute median $m^{2}=||x-x^{\prime}||^{2}_{2}$ for (rbf) kernel and $m=||x-x^{\prime}||_{2}$ for (exp) kernel, on a subset of $100$ batches of size $512$ from the dataset. Then, for (rbf) kernel, we define $\sigma^{2}=m^{2}\gamma$ and we sweep over $\gamma\in\{0.01,0.1,0.3,0.5,1.0,1.5,2.0,10.0\}$. For ($\exp$) kernel, we define $\sigma=m\gamma$ and we sweep over $\gamma\in\{0.01,0.1,0.3,0.5,1.0,1.5,2.0,10.0\}$. Since we only use kernel diffusion losses on CIFAR-10, we only report the value of the medians in this case. We have that $m^{2}=1357.8127$ and $m=36.9$.
@@ I.3 Robotics experiments
We used the Libero benchmark (Liu et al., 2024), a life-long learning benchmark that consists of 130 language-conditioned robotic manipulation tasks. There are 5 suites in the Libero benchmark, Libero-Spatial, Libero-Object, Libero-Goal, Libero-Long and Libero-90. Each suite contains 10 tasks, except Libero-90 which has 90 tasks. In our experiments we focus on Libero-Long, the most challenging suite with 10 tasks that features long-horizon tasks with diverse object interactions. We train language-conditioned multitask diffusion policies for the 10 tasks of Libero10 suite. and evaluate the success rates of the policies in simulation. We encode the visual observations using a ResNet (He et al., 2016), and encode the language instruction using a pretrained and frozen Bert (Devlin et al., 2018) encoder. Our multitask diffusion policy does action chunking (Chi et al., 2023), and predicts and executes a chunk of 8 7-dimensional actions conditioned on the language instruction, the current visual observations and proprioceptive states. For the denoiser, we follow Aloha unleashed (Zhao et al., 2024) and use a cross-attention based transformer denoiser.
We follow the original Libero paper for training and evaluating our multitask diffusion policy. As a loss function for training the diffusion policy, we use the Mean Absolute Error(MAE), i.e., $\ell_{1}$ loss. See Section J.2 for results with Mean Squared Error (MSE). Libero10 suite has 138090 steps of training data and we use a batch size of 32 to train our policies for 250k steps (roughly 50 epochs). Evaluation is on simulated environment for 50 episodes per task following the Libero benchmark. We evaluate our policy every 25k steps (roughly 5 epochs). We take the median success rate of the policy across all tasks in the suite and repeat the experiment over 3 seeds. We report the median success rates over all tasks and best checkpoint during training.
@@
Appendix J Additional experiments
J.1 Additional image experiments
@@
J.2 Additional robotics experiments
@@
In the setting of Section 6.3, we only consider the case $\beta=1$ and compared our results for several values of $\lambda\in[0,1]$, namely $\lambda\in\{0.0,0.1,0.5\}$. In Figure 19, we present a similar study as Figure 9 but in the case where $\beta=2$. Note that the diffusion baseline is not the same in Figure 19 and in Figure 9. In Figure 9, the diffusion baseline is trained with the Mean Absolute Error (MAE) regression loss, i.e., we replace the $\ell_{2}^{2}$ loss by the $\ell_{1}$ loss in the diffusion model loss. However, in Figure 19, the diffusion model baseline is trained with the classical MSE regression loss.
@@
J.3 Additional samples
For CIFAR-10, samples from the distributional model $\hat{x}_{\theta}$ trained with $\beta=0.1,\lambda=1$ and with $10$ sampling steps are shown in Figure 20, left. Samples from the one trained with $\beta=2,\lambda=0$ and with $100$ sampling steps are shown in Figure 21, leeft. The samples from the diffusion model with $10$ steps are shown in Figure 20, right and with $100$ steps are shown in Figure 21, right.

### v3

@@ F.2 Connection with DDIM
Note that in (Song et al., 2021a), $\alpha_{t}^{2}$ is replaced with $\alpha_{t}$ and $\eta_{s,t}$ is denoted $\sigma_{t}$. Recall that in our introduction of diffusion models in Section 2.1, see (6) we have
$\mu_{s,t}(x_{0},x_{t})=(\varepsilon^{2}r_{1,2}(s,t)+(1-\varepsilon^{2})r_{0,1})x_{t}$ (174)
@@ Appendix H Extended related work
Relationship to works using $\mathrm{MMD}$.
Scoring rules have corresponding $\mathrm{MMD}$ divergences, and approaches exist that use the $\mathrm{MMD}$ in diffusion and particle flow models. Galashov et al. (2025) generates a sequence of distributions from a forward diffusion process. It then generates noise-dependent neural $\mathrm{MMD}$s between clean and noisy data, and performs gradient flow to move particles from noise level $t$ to a lower noise level $s<t$. Compared to ours, Galashov et al. (2025) does not use a generator since it is a particle flow. Their best CIFAR-10 FID is $7.7$. Aiello et al. (2024) first trains a diffusion model, and then refines/distills by coarsening the reverse timesteps, and uses MMD on CLIP features to finetune the DDM denoiser. Their CIFAR10 FID for $NFE=10$ improves from $13.6$ to $3.8$, while ours is $3.19$.
@@ I.2 Image space experiments
For all the models, we sweep over weighting $w_{t}$ (see (8), (14)), where we either do not use any weighting or we use sigmoid weighting, see (Kingma et al., 2021), weighting with bias parameters $\{-2,-1,0,1,2\}$. When we train classical diffusion model, we always use velocity as a prediction target, i.e., the target is $X_{0}-X_{1}$. When we train distributional model, we use $X_{0}$ as prediction target. For distributional models, we additionally sweep over $\lambda\in\{0,0.1,0.5,1.0\}$, $\beta\in\{0.0001,0.001,0.01,0.1,0.5,1.0,1.5,1.99,2.0\}$. We use population size $m=4$. We also tried $m=2$ but it led to slightly worse results (the resulting FID was $7\%$ worse). We did not use higher population sizes since it led to higher computational and memory cost (see Appendix J for more details). When we use kernel diffusion losses, see  Section 2.2 and Section 3, for (imq) kernel (9a), we sweep over $c\in\{0.01,0.1,0.3,0.5,1.0,1.5,2.0,10.0,50.0,100.0\}$. For for (rbf) kernel (9b) and for (exp) kernel (9c), we firstly compute median $M^{2}=||x-x^{\prime}||^{2}_{2}$ for (rbf) kernel and $M=||x-x^{\prime}||_{2}$ for (exp) kernel, on a subset of $100$ batches of size $512$ from the dataset. Then, for (rbf) kernel, we define $\sigma^{2}=M^{2}\gamma$ and we sweep over $\gamma\in\{0.01,0.1,0.3,0.5,1.0,1.5,2.0,10.0\}$. For ($\exp$) kernel, we define $\sigma=M\gamma$ and we sweep over $\gamma\in\{0.01,0.1,0.3,0.5,1.0,1.5,2.0,10.0\}$. Since we only use kernel diffusion losses on CIFAR-10, we only report the value of the medians in this case. We have that $M^{2}=1357.8127$ and $M=36.9$.
@@ I.3 Robotics experiments
We used the Libero benchmark (Liu et al., 2024), a life-long learning benchmark that consists of 130 language-conditioned robotic manipulation tasks. There are 5 suites in the Libero benchmark, Libero-Spatial, Libero-Object, Libero-Goal, Libero-Long and Libero-90. Each suite contains 10 tasks, except Libero-90 which has 90 tasks. In our experiments we focus on Libero-Long, the most challenging suite with 10 tasks that features long-horizon tasks with diverse object interactions, and reproduce the main experiment settings on three other suites: Libero-Spatial, Libero-Object, Libero-Goal, with 10 tasks. We train language-conditioned multitask diffusion policies for the 10 tasks of Libero10 suite and evaluate the success rates of the policies in simulation. We encode the visual observations using a ResNet (He et al., 2016), and encode the language instruction using a pretrained and frozen Bert (Devlin et al., 2018) encoder. Our multitask diffusion policy does action chunking (Chi et al., 2023), and predicts and executes a chunk of 8 7-dimensional actions conditioned on the language instruction, the current visual observations and proprioceptive states. For the denoiser, we follow Aloha unleashed (Zhao et al., 2024) and use a cross-attention based transformer denoiser.
We follow the original Libero paper for training and evaluating our multitask diffusion policy. As a loss function for training the diffusion policy, we use the Mean Squared Error (MSE). Libero10 suite has 138090 steps of training data and we use a batch size of 32 to train our policies for 250k steps (roughly 50 epochs). Evaluation is on simulated environment for 50 episodes per task following the Libero benchmark. We evaluate our policy every 25k steps (roughly 5 epochs). We take the median success rate of the policy across all tasks in the suite and repeat the experiment over 3 seeds. We report the median success rates over all tasks and best checkpoint during training. We use the same set for other three suites and train for 50 epochs and evaluate every 5 epochs.
@@
Appendix J Computational complexity
Our method has the same computational complexity during sampling as ordinary diffusion models. However, it has an increased computational complexity during training which is detailed below.
Assume that the computational complexity of the forward pass in our diffusion model is $O(F)$ and that the dimensionality of $x$ is $D$.
Diffusion loss case. The loss function in (15) with $m=1$ and $\lambda=0$, can be thought of as a standard diffusion model loss function. It requires $O(nF)$ evaluations to compute $x_{\theta}(t_{i},X_{t_{i}})$ for every element of a batch. Then, to evaluate the loss (the norm), the complexity is $O(nD)$. Therefore the total cost of computing the loss is $O(n(F+D))$. The backwards pass is proportional to the forward pass and has computational complexity $O(nF)$.
Distribution loss case. For $\lambda\neq 0$ and $m>1$, we first need $O(nmF)$ function evaluations to compute $x_{\theta}(t_{i},X_{t_{i}},\xi_{i,j})$ for $i=1,\ldots,n$ and for $\xi_{i,j}$, $j=1,\ldots,m$. Additionally, we need $O(nmD)$ operations to compute the first (diffusion-like) terms of the loss. Then, in order to compute the interaction terms, we need $O(nm(m-1)D)$ time. Therefore, the total cost is of computing the loss is $O(nm(F+D)+nm(m-1)D)$. Naively, the backwards pass will take $O(nmF+nm(m-1)F)=O(nm^{2}F)$ time. However, since the 2nd term uses the gradients of $x_{\theta}(t_{i},X_{t_{i}},\xi_{i,j})$, one could precompute these for all $i=1,\ldots,n$ and $j=1,\ldots,m$ and therefore decrease the total backwards cost to $O(nmF)$. We found that in practice, the XLA compiler in JAX performs this optimization without explicit coding, see our results below.
Runs on real hardware. We compared the training times on real hardware. We study the training of CIFAR-10 as we vary $m$. We report steps per second metric where a step corresponds to a full forward+backward step. As hardware, we use A100 GPU (40Gb of memory) with batch size = 16, H100 GPU (80Gb of memory) with batch size = 64 and TPUv5p (95Gb of memory) with batch size = 64 (per device with 4 devices in total). The results below indicate that steps per second decreases proportionally to $m$.
Hardware Diffusion Distributional (m=2) Distributional (m=4) Distributional (m=8)
A100 (batch size =16) 9.05 6.78 4.6 2.77
H100 (batch size =64) 14.3 7.85 4.15 2.06
TPUv5 (batch size =64) 11.2 8.5 4.3 2.22
Table 1: Steps per seconds on real hardware for different models.
Appendix K Additional experiments
K.1 Additional 2D experiments
We trained an unconditional standard diffusion model and our distributional models on a more complex 2D distribution – checkerboard, see Galashov et al. (2025). We report MMD between sampled and target distributions with different NFEs, smaller MMD is in bold, in Table 2 Our approach outperforms standard diffusion for small NFEs.
NFEs Diffusion Distributional
5 7.00e-3 1.57e-4
10 2.10e-3 1.96e-4
50 2.70e-4 1.55e-4
100 1.96e-4 2.00e-4
1000 1.71e-4 1.83e-4
Table 2: MMD between target and sampled distribution for checkerboard dataset.
K.2 Additional image experiments
@@
ImageNet.
We ran initial experiments on ImageNet (Russakovsky et al., 2015) with resolution $64\times 64\times 3$, but so far were unable to obtain satisfactory results. We will focus on scaling our method up to large datasets in the follow-up work.
K.3 Comparisons to distillation and numerical solvers
In this section, we compare our approach with the DPM-solver++ method Lu et al. (2022) and the state-of-the-art, multi-step distillation method, moment-matching distillation (Salimans et al., 2024).
For all the sampling methods, we sweep over the “safety” parameter $\eta$, which specifies the time interval to be $[\eta,1-\eta]$. We considered values $\{0.1,0.01,0.001,0.0001\}$. Moreover, for all the methods we also sweep over the “churn” parameter $\varepsilon$ which controls the stochasticity, see Equation 6, we tried $\{0.0,0.25,0.5,0.75,1.0\}$. We found the stochastic version of DPM-solver++ (with a churn parameter close to $1$) overall worked the best. For ordinary diffusion and our method, we use a “uniform” time schedule. For DPM-solver++ we use the “logsnr” time schedule (see Lu et al. (2022)), which led to the best results (we also tried EDM and uniform).
For our distillation comparisons, we distill pre-trained diffusion models into student models. These student models are trained for 100,000 iterations, using a batch size of 256 for pixel-space models and 32 for latent-space models. During training, we sweep over the learning rate and Exponential Moving Average (EMA) decay.
The student model utilizes a stochastic DDIM sampler with a churn parameter, $\varepsilon$. We sweep over $\varepsilon$ values in the set $\{0.0,0.5,1.0\}$. We found that at sampling time, setting the churn parameter to $\varepsilon=0$ for the distilled student yielded the best results.
For all the methods, we present the results with the best hyperparameters in the tables below (Table 3,Table 4,Table 5,Table 6).
Table 3: CIFAR-10 (Conditional) results. The metric is FID averaged across 3 random seeds. We report results with different NFEs for baseline diffusion model, DPM++, Distributional (ours) and moment matching distillation. In bold, we highlight the method with the lowest FID and in italic, we highlight the method with the second lowest FID.
NFEs Diffusion DPM++ Distributional Distillation
2 80.77 19.41 29.33 5.19
4 23.31 12.60 4.67 3.84
8 7.53 3.95 3.21 3.13
10 5.67 3.55 3.19 2.99
12 4.83 3.69 3.03 2.93
15 3.61 3.91 2.87 2.76
Table 4: CelebA (Conditional) results. The metric is FID averaged across 3 random seeds. We report results with different NFEs for baseline diffusion model, DPM++, Distributional (ours) and moment matching distillation. In bold, we highlight the method with the lowest FID and in italic, we highlight the method with the second lowest FID.
NFEs Diffusion DPM++ Distributional Distillation
2 62.85 58.39 81.79 7.02
4 26.25 20.39 22.98 6.47
8 13.07 17.28 4.94 4.43
10 11.67 22.48 3.60 4.43
12 10.00 23.23 3.37 4.25
15 8.51 23.53 3.45 4.27
Table 5: LSUN (Unconditional) results. The metric is FID averaged across 3 random seeds. We report results with different NFEs for baseline diffusion model, DPM++, Distributional (ours) and moment matching distillation. In bold, we highlight the method with the lowest FID and in italic, we highlight the method with the second lowest FID.
NFEs Diffusion DPM++ Distributional Distillation
2 238.85 120.99 233.24 16.04
4 41.57 43.60 69.90 5.54
8 10.82 12.95 9.79 4.52
10 9.08 17.85 7.14 4.64
12 7.04 16.07 6.12 3.81
15 6.51 23.91 5.78 4.39
Table 6: Latent CelebA-HQ (Unconditional) results. The metric is FID averaged across 3 random seeds. We report results with different NFEs for baseline diffusion model, DPM++, Distributional (ours) and moment matching distillation. In bold, we highlight the method with the lowest FID and in italic, we highlight the method with the second lowest FID.
NFEs Diffusion DPM++ Distributional Distillation
2 101.47 114.44 67.60 37.88
4 53.83 26.05 23.80 14.47
8 15.60 20.70 9.84 8.22
10 11.89 21.84 8.15 6.76
12 9.97 24.61 7.04 6.05
15 8.39 27.31 6.21 5.46
The results indicate that our approach is competitive with multistep distillation (across a varied number of steps) and outperforms DPM-Solver++ at 8 or more Number of Function Evaluations (NFEs). The performance of DPM-Solver++ degrades as the number of steps increases, which is an expected consequence of its numerical instability.
Our approach has two compelling advantages over multistep distillation. First, it does not rely on distillation, eliminating the need to train a large, slow teacher model. Second, it does not require specifying additional sampler hyperparameters during training.
Furthermore, combining our distributional approach with a modern distillation method is a potentially interesting topic for future research.
K.4 Additional robotics experiments
In addition to Libero10 suite, we repeated our main experiment on three other suites in Libero benchmark: Liber-Goal, Libero-Spatial, and Libero-Object. We present the performance on these three additional suites in Figure 19. In all three suites, distributional diffusion helps the performance when fewer number of diffusion steps are used, reproducing our results on Libero10.
@@
In the setting of Section 6.3, we only consider the case $\beta=1$ and compared our results for several values of $\lambda\in[0,1]$, namely $\lambda\in\{0.0,0.1,0.5\}$. In Figure 20, we present a similar study as Figure 9 but in the case where $\beta=2$.
We also ran an ablation over $m$ with fixed $\beta=1$ and $\lambda=0.5$. The results are reported in Table 7.
$m$ Success rate
2 0.783
4 0.73
8 0.817
16 0.800
32 0.840
64 0.797
Table 7: Ablation over $m$ for robotics dataset.
Figure 19: Performance in three other Libero suites - Libero-Goal, Libero-Spatial, and Libero-Object - as function of diffusion steps and $\lambda$ with same settings as Figure 9. We plot the maximum success rate from best checkpoint during training. Distributional models consistently performed the best when fewest diffusion steps are used. $\lambda>0$ also performs better than $\lambda=0$
@@
K.5 Additional samples
For CIFAR-10, samples from the distributional model $\hat{x}_{\theta}$ trained with $\beta=0.1,\lambda=1$ and with $10$ sampling steps are shown in Figure 21, left. Samples from the one trained with $\beta=2,\lambda=0$ and with $100$ sampling steps are shown in Figure 22, left. The samples from the diffusion model with $10$ steps are shown in Figure 21, right and with $100$ steps are shown in Figure 22, right.